The second of two numbers is 3 more than the first. If 4 times the first is increased by the second, the result is 73. Find the numbers.
Answer by funmath(2933)

Let the 1st number=x
Then the 2nd number=x+3
4 times (4*) the first number (x) increased by (+) the second number (x+3)the result is (=) 73.
4(x)+(x+3)=73
4x+x+3=73
5x+3=73
5x+3-3=73-3
5x=70
5x/5=70/5
x=14
The first number:x=14
The second number:x+3=14+3=17
Happy calculating!!!